40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3929     Entry time: Tue Nov 16 03:33:22 2010
Author: yuta 
Type: Update 
Category: IOO 
Subject: aligned Faraday, beam reached SM just before PRM 

(Koji, Yuta)

We aligned the Faraday after MC and we are now ready to install PRM.

Background:
  MC was roughly aligned (beam spot ~0.7mm from the actuation center).
  So, we started aligning in-vac optics.
  First thing to align was the Faraday after MC3.

What we did:
  1. Ran A2L.py for confirmation.(Second from the last measurement point on the A2L result plot)

  2. Aligned the Faraday so that MC3 trans can go through it. We moved the Faraday itself, while we didn't touch IM2.
     We turned the pitch nob of the last steering mirror at PSL table in CCW slightly in order to lower the beam at the Faraday by ~1mm.

  3. During the alignment, we found that the polarization of the incident beam was wrong. It should have been S but it was P.
     As there is the HWP right before the EOM, Rana rotated it so as to have the correct polarization of S on the EOM and the MC.
     Note that the PMC and the main interferometer are configured to have P-pol while the MC is to have S-pol.

  4. Setup the video camera to monitor the entrance aperture of the Faraday. It required 4 steering mirrors to convey the image to the CCD.

  5. Moved all of the OSEMs for MC1 and MC3 so that the sensor output can have roughly half of their maxima.

  6. Ran A2L.py. (The last measurement point on the A2L result plot)

  7. Aligned the IO optics so that the beam goes Faraday -> MMT1 -> MMT2 -> SM3.

Result:

  1. OSEM sensor outputs for MC1 and MC3 are;

(V) MC1 MC3
max current value max current value
ULSEN 1.3 0.708 1.37 0.699
URSEN 1.4 0.845 1.71 0.796
LRSEN 1.45 0.743 1.77 0.640
LLSEN 1.56 0.762 1.56 0.650
SDSEN 1.67 0.801 1.59 0.821



  2. A2L result is;
MCalignNov16.png


     The beam position slightly got lower(~0.2mm), because we touched SM at PSL table.
     Alignment slider values changed because we moved MC1 and MC3 OSEMs.

  3. Now, MC_RFPD_DCMON is ~0.39 when MC unlocked and ~0.083 when locked.
     So, the visibility of MC is ~79% (for S-pol).

  4. Now the incident beam to the MC has S polarization, the cavity has higher finesse. This results the increased MC trans power.
     It was ~8e2 when the polarization was P, now it is ~4.2e3 when the MC is locked.

  5. The beam reached SM3 at BS table. The alignment of the SM2, MMT1, MMT2 were confirmed and adjusted.

  6. All pieces of the leftover pizza reached my stomach.

Plan:
  - Install PRM to the BS chamber.
  - Align PRM and get IFO reflection beam out to the AP table
 

ELOG V3.1.3-